jhonyy9 (jhonyy9):

- how and when it is possible or we need using the reductio ad absurdum prove-style ?- with examples please -

OpenStudy (anonymous):

a classic example is to prove that \[\sqrt{2}\] is irrational. this is because you are trying to prove it is NOT something. obviously you do not have time to check every fraction, so you ASSUME it is a fraction and then use "reductio ad absurdum" to show that you end up with something false. this means your assumption was not true, i.e. in this case that \[\sqrt{2} \] is NOT rational

jhonyy9 (jhonyy9):

ok ty. but for one equation like m=a+b for example - how us ? please

OpenStudy (anonymous):

i am not sure what you mean \[m=a+b\] is an equation, not a theorem. so you cannot "prove" it by any method

OpenStudy (anonymous):

a theorem comes on the form of " if something is true, then we know something else must also be true"

jhonyy9 (jhonyy9):

for example if we want to prove it that m is always even if a and b are odd

OpenStudy (anonymous):

well that i think you can prove directly. i would not use it in this case. it is silly to use a proof by contradiction when you can prove something directly because it keeps you from explaining why exactly it works

OpenStudy (anonymous):

this take three lines 1) if a and b are odd then \[a=2j+1,b=2k+1\] 2)\[a+b=2k+1+2j+1=2(j+k+1)\] is even

OpenStudy (anonymous):

ok two lines

jhonyy9 (jhonyy9):

ok and if m is even can be write always in the form 2n - so than if a+b=2(j+k+1)=2n than n=j+k+1 yes ?

OpenStudy (anonymous):

"The opposite of every great idea is another great idea." "This statement is false"